Will observe J1713+0747 for 900 seconds in both search and fold modes, and B1442+101 as a fluxcal for 90 seconds
Session begins at 17:00 EDT.
Boom extended, new vegas.conf copied into /home/gbt/etc/config with new BOF files
First scan is on B1442+101 at 350 MHz with i0100x8192 mode
Levels were a low with vegas.scale set to 200. Took more pairs of scans with vegas.scale = 800, 400, 300, and 250 but all were too high. Finally settled on back on vegas.scale = 200. Maybe balance was a bit off last time.
Use scans 11/12
J1713+0747 is not up (I didn't check the LST I was given, but I guess it wasn't what I asked for). Using J1125+7819 instead.
All observations went smoothly
Switched to L-Band to test c0800x0064 mode
Won't have time for fluxcal scan in this mode
All seems to have gone well. Leaving these bof files in place.